OpenStudy (anonymous):

Express the complex number in trigonometric form. 3 - 3i

OpenStudy (anonymous):



OpenStudy (anonymous):

you need \(r\) and \(\theta\) \(r=\sqrt{a^2+b^2}=\sqrt{3^2+3^2}=3\sqrt{2}\)

OpenStudy (anonymous):

you should see from the picture i drew that \(\theta=-\frac{\pi}{4}\)

OpenStudy (anonymous):

oh not a choice, i see, ok you can use \(\theta=\frac{7\pi}{4}\) same place on the unit circle

OpenStudy (anonymous):

is it the third choice

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it